Three out of 5 people attending Dustin's party were girls. If 12 girls attended, how many people were at the party?
Answer by wuwei96815(245)

Well, we don't know how many people attended so let's call that number x the unknown.

Now, we have to write an equation with x in it so that we can solve the equation and find out what x is.

3 out of 5 people attending or 3/5 of x were 12 girls.
Another way of saying the same thing is 3/5 x = 12.
Let's solve 3/5x=12.
Multiply both sides of the equation by 5/3 in order to isolate x.
5/3 times 3/5x = 5/3 times 12.
We get: x = 20.

Did you understand all of that?